OpenStudy (anonymous):

HELPPPPPPPPPPPPPPPPP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! sometime soon please.... 24. Calculate a. a. 6 b. 10 c. 2.67 d. 13.5

OpenStudy (anonymous):



Directrix (directrix):

Write here the theorem that begins "The bisector of an angle of a triangle divides the opposite side into ___________________?_______________. Once you do that, somebody will help you solve the problem.

OpenStudy (anonymous):

k An angle bisector of an angle of a triangle divides the opposite side in two segments that are proportional to the other two sides of the triangle.

Directrix (directrix):

That means the following: 9 is to 6 as a is to 4. 9 / 6 = a / 4 6a = 9*4 6a = 36 a = 6

Directrix (directrix):

Question?

OpenStudy (anonymous):

thank you i understand now